soy bean meal is 12% protein cornmeal is 6% protein How many pounds of each should be mixed together in order to get 240 pound mixture that is 7% protein?
How many pounds of the cornmeal should be in the mixture?
How many pounds of soybean meal should be in the mixture?

Answer by orca(409)

Let x be the amount of cornmeal to be mixed.
Then the amount of soy bean meal to be mixed is 240 - x.
The amount of protein in the cornmeal is x*6%.
The amount of protein in the soy bean meal is (240 - x)*12%.
The amount of protein in the 240 pounds mixture is 240*7%.
As the amount of protein in the cornmeal + the amount of protein in the soy bean meal = amount of protein in the 240 pounds mixture, we have:
x*6% + (240 - x)*12% = 240*7%
Solving the equation for x, we have
6x + 12(240 - x) = 7*240 (multiplying both sides by 100)
6x + 2880 - 12x = 1680
-6x = -1200
x = 200
So 200 pounds of the cornmeal should be in the mixture.